# Computing Exact Shapley Values in Polynomial Time for Product-Kernel Methods

### v1

@@ C.3 Proof of Theorem 4
C.4 Proof of Lemma 5
By the efficiency property of Shapley values [1], the sum of Shapley values equals the difference between the value of the grand coalition and the empty coalition:
@@ C.4 Proof of Lemma 5
From Proposition 3, we have:
@@
C.5 Proof of Theorem 6
  • (i)
    We begin by expressing the full kernel function as a product of feature-wise kernels:
    $k(\boldsymbol{x}^{(i)},\boldsymbol{x}^{(j)})=\prod_{q\in\mathcal{D}}k_{q}(x^{(i)}_{q},x^{(j)}_{q}).$
    Using Theorem 10, we expand the product:
    $\prod_{q\in\mathcal{D}}k_{q}(x^{(i)}_{q},x^{(j)}_{q})=\sum_{\mathcal{S}\subseteq\mathcal{D}}\prod_{q\in\mathcal{S}}\big{(}k_{q}(x_{q}^{(i)},x_{q}^{(j)})-1\big{)}.$
    Substituting this into the MMD formulation for $k(\boldsymbol{x}^{(i)},\boldsymbol{x}^{(j)})$, $k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})$, and $k(\boldsymbol{x}^{(i)},\boldsymbol{z}^{(j)})$ yields
    $\widehat{\text{MMD}}^{2}(\mathbb{P},\mathbb{Q})=\frac{1}{n(n-1)}\sum_{i\neq j}\sum_{\mathcal{S}\subseteq\mathcal{D}}\prod_{q\in\mathcal{S}}\big{(}k_{q}(x^{(i)}_{q},x^{(j)}_{q})-1\big{)}$
    $\quad+\frac{1}{m(m-1)}\sum_{i\neq j}\sum_{\mathcal{S}\subseteq\mathcal{D}}\prod_{q\in\mathcal{S}}\big{(}k_{q}(z^{(i)}_{q},z^{(j)}_{q})-1\big{)}-\frac{2}{nm}\sum_{i,j}\sum_{\mathcal{S}\subseteq\mathcal{D}}\prod_{q\in\mathcal{S}}\big{(}k_{q}(x^{(i)}_{q},z^{(j)}_{q})-1\big{)}.$
    By the linearity of summation, we have
    $\widehat{\text{MMD}}^{2}(\mathbb{P},\mathbb{Q})$ $=\sum_{\mathcal{S}\subseteq\mathcal{D}}\bigg{[}\frac{1}{n(n-1)}\sum_{i\neq j}\prod_{q\in\mathcal{S}}\big{(}k_{q}(x^{(i)}_{q},x^{(j)}_{q})-1\big{)}$
    $\quad+\frac{1}{m(m-1)}\sum_{i\neq j}\prod_{q\in\mathcal{S}}\big{(}k_{q}(z^{(i)}_{q},z^{(j)}_{q})-1\big{)}-\frac{2}{nm}\sum_{i,j}\prod_{q\in\mathcal{S}}\big{(}k_{q}(x^{(i)}_{q},z^{(j)}_{q})-1\big{)}\bigg{]}.$
    This establishes the functional decomposition of MMD:
    $\widehat{\text{MMD}}^{2}(\mathbb{P},\mathbb{Q})=\sum_{\mathcal{S}\subseteq\mathcal{D}}f^{\text{MMD}}_{\mathcal{S}}(\mathbb{P},\mathbb{Q}),$
    where $f^{\text{MMD}}_{\mathcal{S}}$ is defined as
    $f^{\text{MMD}}_{\mathcal{S}}(\mathbb{P},\mathbb{Q})$ $=\frac{1}{n(n-1)}\sum_{i\neq j}\prod_{q\in\mathcal{S}}\big{(}k_{q}(x^{(i)}_{q},x^{(j)}_{q})-1\big{)}$
    $+\frac{1}{m(m-1)}\sum_{i\neq j}\prod_{q\in\mathcal{S}}\big{(}k_{q}(z^{(i)}_{q},z^{(j)}_{q})-1\big{)}-\frac{2}{nm}\sum_{i,j}\prod_{q\in\mathcal{S}}\big{(}k_{q}(x^{(i)}_{q},z^{(j)}_{q})-1\big{)}.$
  • (ii)
    Following a similar approach as in Proposition 3, we utilize the property that if a function admits a functional decomposition, the corresponding value function for the cooperative game is uniquely defined as:
    $v_{\text{MMD}}(\mathcal{S})=\sum_{\mathcal{T}\subseteq\mathcal{S}}f^{\text{MMD}}_{\mathcal{T}}(\mathbb{P},\mathbb{Q}).$
    Applying this to MMD, we define:
    $v_{\text{MMD}}(\mathcal{S})$ $=\sum_{\mathcal{T}\subseteq\mathcal{S}}\frac{1}{n(n-1)}\sum_{i\neq j}\prod_{q\in\mathcal{S}}\big{(}k_{q}(x^{(i)}_{q},x^{(j)}_{q})-1\big{)}$
    $\qquad+\frac{1}{m(m-1)}\sum_{i\neq j}\prod_{q\in\mathcal{S}}\big{(}k_{q}(z^{(i)}_{q},z^{(j)}_{q})-1\big{)}-\frac{2}{nm}\sum_{i,j}\prod_{q\in\mathcal{S}}\big{(}k_{q}(x^{(i)}_{q},z^{(j)}_{q})-1\big{)}$
    $=\frac{1}{n(n-1)}\sum_{i\neq j}\sum_{\mathcal{T}\subseteq\mathcal{S}}\prod_{q\in\mathcal{S}}\big{(}k_{q}(x^{(i)}_{q},x^{(j)}_{q})-1\big{)}$
    $+\frac{1}{m(m-1)}\sum_{i\neq j}\sum_{\mathcal{T}\subseteq\mathcal{S}}\prod_{q\in\mathcal{S}}\big{(}k_{q}(z^{(i)}_{q},z^{(j)}_{q})-1\big{)}-\frac{2}{nm}\sum_{i,j}\sum_{\mathcal{T}\subseteq\mathcal{S}}\prod_{q\in\mathcal{S}}\big{(}k_{q}(x^{(i)}_{q},z^{(j)}_{q})-1\big{)}.$
    By applying Theorem 10 to the inner sum, we can simplify the above equation as:
    $v_{\text{MMD}}(\mathcal{S})=\frac{1}{n(n-1)}\sum_{i\neq j}k_{\mathcal{S}}(\boldsymbol{x}^{(i)}_{\mathcal{S}},\boldsymbol{x}^{(j)}_{\mathcal{S}})+\frac{1}{m(m-1)}\sum_{i\neq j}k_{\mathcal{S}}(\boldsymbol{z}^{(i)}_{\mathcal{S}},\boldsymbol{z}^{(j)}_{\mathcal{S}})-\frac{2}{nm}\sum_{i,j}k_{\mathcal{S}}(\boldsymbol{x}^{(i)}_{\mathcal{S}},\boldsymbol{z}^{(j)}_{\mathcal{S}}),$
    which completes the proof. ∎
C.6 Proof of Proposition 7
@@
C.7 Proof of Theorem 8
  • (i)
    We begin by expressing the full kernel matrix as a product of feature-wise kernels:
    $\mathbf{K}=\bigodot_{j\in\mathcal{D}}\mathbf{K}_{j}.$
    since $\mathbf{K}$ is element-wise products of $\mathbf{K}_{j}$, we can apply Theorem 10 for each element in the kernel matrix and write:
    $\bigodot_{j\in\mathcal{D}}\mathbf{K}_{j}=\sum_{\mathcal{S}\subseteq\mathcal{D}}\bigodot_{j\in\mathcal{S}}(\mathbf{K}_{j}-\boldsymbol{1}\boldsymbol{1}^{\top}).$
    Since HSIC is dependent on $\mathrm{tr}(\mathbf{K}\mathbf{H}\mathbf{L}\mathbf{H})$, substituting the decomposition of $\mathbf{K}$ gives:
    $\widehat{\text{HSIC}}(X,y)=\frac{1}{(n-1)^{2}}\mathrm{tr}\left(\mathbf{H}\mathbf{L}\mathbf{H}\sum_{\mathcal{S}\subseteq\mathcal{D}}\bigodot_{j\in\mathcal{S}}(\mathbf{K}_{j}-\boldsymbol{1}\boldsymbol{1}^{\top})\right).$
    Using the linearity of the trace operator, we obtain:
    $\widehat{\text{HSIC}}(X,y)=\frac{1}{(n-1)^{2}}\sum_{\mathcal{S}\subseteq\mathcal{D}}\mathrm{tr}\left(\mathbf{H}\mathbf{L}\mathbf{H}\bigodot_{j\in\mathcal{S}}(\mathbf{K}_{j}-\boldsymbol{1}\boldsymbol{1}^{\top})\right).$
    This completes the proof of the functional decomposition.
  • (ii)
    Similar to Proposition 3, we define the value function for a cooperative game as
    $v_{\text{HSIC}}(\mathcal{S})=\sum_{\mathcal{T}\subseteq\mathcal{S}}\frac{1}{(n-1)^{2}}\mathrm{tr}\left(\mathbf{H}\mathbf{L}\mathbf{H}\bigodot_{j\in\mathcal{T}}(\mathbf{K}_{j}-\boldsymbol{1}\boldsymbol{1}^{\top})\right).$
    By applying Theorem 10, we simplify:
    $\sum_{\mathcal{T}\subseteq\mathcal{S}}\bigodot_{j\in\mathcal{T}}(\mathbf{K}_{j}-\boldsymbol{1}\boldsymbol{1}^{\top})=\bigodot_{j\in\mathcal{S}}\mathbf{K}_{j}.$
    Thus, the value function for feature subset $\mathcal{S}$ is:
    $v_{\text{HSIC}}(\mathcal{S})=\frac{1}{(n-1)^{2}}\mathrm{tr}\left(\mathbf{H}\mathbf{L}\mathbf{H}\bigodot_{j\in\mathcal{S}}\mathbf{K}_{j}\right)=\frac{1}{(n-1)^{2}}\mathrm{tr}\left(\mathbf{H}\mathbf{L}\mathbf{H}\mathbf{K}_{\mathcal{S}}\right),$
    which completes the proof. ∎
C.8 Proof of Proposition 9
@@
Appendix D Recursive and Numerically Stable Algorithms for Computing Shapley Values for Product-Kernel Learning Models
@@ D.1 Recursive Algorithm
Though being efficient, the recursive formulation suffers from numerical instability due to alternating sign cancellation, power operations, and error amplification through division. To develop a stable approach, we use the fact that symmetric elementary polynomials emerge naturally as coefficients of the characteristic polynomial [34]:
@@
[…]
11:        Append $-\tilde{\boldsymbol{z}}_{i}\cdot\text{Coeff}[0]$ to new_coeff {Term for $\boldsymbol{x}^{0}$}
@@
Appendix E Additivity of Explanations for Learning Models, MMD and HSIC
E.1 Explanation Additivity for Learning Models
@@ E.2 Explanation Additivity for MMD
By the efficiency property of Shapley values [1]:
@@
E.3 Explanation Additivity for HSIC
@@
By the efficiency property of Shapley values [1], the sum of Shapley values equals the difference between the value of the grand coalition and the empty coalition:
@@
From Theorem 8 (ii), we have:
@@
Appendix F HSIC attribution with two multivariate variables
@@
    We first assume that $\mathbf{L}$ is fixed, and try to attribute the total HSIC to the variables in $X$. Since $k$ is a product kernel, we can recover a functional decomposition based on $k$, similar to Theorem 8. Let $\mathcal{D}_{X}$ be the set of variables of $X$, we therefore define the value function for attributing the total HSIC to $X$ variables as:
@@
    By the same token, we take $\mathbf{K}$ fixed and try to attribute the total HSIC to the variables in $Z$. Since $l$ is a product kernel, we can recover a functional decomposition based on $l$ and define the value function as:
@@
Appendix G Experiments
G.1 Experimental Setup
SVM Optimization Using Optuna [45]
When using SVM in our experiments, we optimized the Support Vector Machine (SVM) with a Radial Basis Function (RBF) kernel using Optuna [45], a robust hyperparameter optimization framework. The target type (either ’regression’ or ’classification’) was determined to guide the selection of the appropriate SVM model (SVR for regression and SVC for classification). The hyperparameters C and gamma, critical for the RBF kernel’s performance, were optimized within an extensive range using a log-uniform distribution. Specifically, we defined the hyperparameters: C between $10^{-5}$ and $10^{5}$ and gamma between $10^{-5}$ and $10^{3}$, and utilized 5-fold cross-validation to ensure reliable evaluation. The optimization process aimed to minimize the mean squared error for regression tasks and maximize accuracy for classification tasks. After conducting the specified number of trials (n=100), the best hyperparameters were used to train a final SVM model on the entire dataset, yielding both the optimal model configuration and the best cross-validation score achieved during the optimization process.
@@ G.1 Experimental Setup
G.2 MMD Experiments
@@
Using MMD, we calculate the dissimilarity between male and female groups and then compute Shapley values to attribute variable contributions to the MMD. Figure 6 displays the Shapley values for the nine variables in the Diabetes dataset. The results show that $s3$ and $s4$ contribute most significantly to the MMD, followed by $bp$, $s6$, $age$, $s5$, and $s2$. In contrast, $bmi$ and $s1$ reduce the MMD, indicating their alignment across the two groups.

### v2

@@ D.1 Proof of Theorem 3
D.2 Proof of Lemma 4
By the efficiency property of Shapley values [Shapley, 1953], the sum of Shapley values equals the difference between the value of the grand coalition and the empty coalition:
@@ D.2 Proof of Lemma 4
For the functional baseline value function, we have:
@@
D.3 Proof of Proposition 6
@@
D.4 Proof of Proposition 8
@@
Appendix E Recursive and Numerically Stable Algorithms for Computing Shapley Values for Product-Kernel Learning Models
@@
Though being efficient, the recursive formulation suffers from numerical instability due to alternating sign cancellation, power operations, and error amplification through division. To develop a stable approach, we use the fact that symmetric elementary polynomials emerge naturally as coefficients of the characteristic polynomial [Egge, 2019]:
@@
10:   $\text{new\_coeff}\leftarrow[\,]$
11:   Append $\tilde{\boldsymbol{z}}_{i}\cdot\text{Coeff}[0]$ to new_coeff {Term for $\boldsymbol{x}^{0}$}
@@
Appendix F Additivity of Explanations for Learning Models, MMD and HSIC
F.1 Explanation Additivity for Learning Models
@@ F.2 Explanation Additivity for MMD
By the efficiency property of Shapley values [Shapley, 1953]:
@@
F.3 Explanation Additivity for HSIC
@@
By the efficiency property of Shapley values [Shapley, 1953], the sum of Shapley values equals the difference between the value of the grand coalition and the empty coalition:
@@
From Definition 7, we have:
@@
Appendix G HSIC attribution with two multivariate variables
@@
    We first assume that $\mathbf{L}$ is fixed, and try to attribute the total HSIC to the variables in $X$. Since $k$ is a product kernel, we can define a value function similar to Definition 7. Let $\mathcal{D}_{X}$ be the set of variables of $X$, we therefore define the value function for attributing the total HSIC to $X$ variables as:
@@
    By the same token, we take $\mathbf{K}$ fixed and try to attribute the total HSIC to the variables in $Z$. Since $l$ is a product kernel, we define the value function as:
@@
Appendix H Experiments
H.1 Experimental Setup
SVM Optimization Using Optuna [Akiba et al., 2019]
When using SVM in our experiments, we optimized the Support Vector Machine (SVM) with a Radial Basis Function (RBF) kernel using Optuna [Akiba et al., 2019], a robust hyperparameter optimization framework. The target type (either ’regression’ or ’classification’) was determined to guide the selection of the appropriate SVM model (SVR for regression and SVC for classification). The hyperparameters C and gamma, critical for the RBF kernel’s performance, were optimized within an extensive range using a log-uniform distribution. Specifically, we defined the hyperparameters: C between $10^{-5}$ and $10^{5}$ and gamma between $10^{-5}$ and $10^{3}$, and utilized 5-fold cross-validation to ensure reliable evaluation. The optimization process aimed to minimize the mean squared error for regression tasks and maximize accuracy for classification tasks. After conducting the specified number of trials (n=100), the best hyperparameters were used to train a final SVM model on the entire dataset, yielding both the optimal model configuration and the best cross-validation score achieved during the optimization process.
@@ H.1 Experimental Setup
H.2 Execution time
To assess the computational efficiency of our recursive algorithm, we conducted a simulation study using a randomly generated kernel function of the form $\boldsymbol{\alpha}^{\top}k(\boldsymbol{x},\mathbf{X})$ with 1000 samples. We computed Shapley values under both the brute-force enumeration and our recursive method by varying the number of features, as plotted in Figure 1. With a runtime budget of 300 seconds, the brute-force computation was feasible only for up to 30 features, already requiring more than 200 seconds at this scale, while exceeding the budget beyond that. In contrast, our recursive algorithm consistently completed the same computations in a fraction of the time, from milliseconds at lower dimensions to under 100 seconds, even with 500 features. These results clearly highlight the substantial efficiency gains and scalability of our method for high-dimensional settings.
H.3 MMD Experiments
@@
Using MMD, we calculate the dissimilarity between male and female groups and then compute Shapley values to attribute variable contributions to the MMD. Figure 7 displays the Shapley values for the nine variables in the Diabetes dataset. The results show that $s3$ and $s4$ contribute most significantly to the MMD, followed by $bp$, $s6$, ageage, $s5$, and $s2$. In contrast, bmibmi and $s1$ reduce the MMD, indicating their alignment across the two groups.
@@
Appendix I How does PKeX-Shapley differ from Mohammadi et al. [2025b]?
In Mohammadi et al. [2025b], the authors propose an exact Shapley computation method for stochastic attribution in FANOVA GP models. Similar to PKeX-Shapley, their approach leverages Newton’s identities and symmetric polynomial representations to obtain closed-form attributions. However, the two methods differ in both scope and capability.
PKeX-Shapley is designed specifically for kernel methods with product kernels. It presents a new value function, i.e., a functional baseline value function, to compute Shapley values. While this formulation enables exact computation of the mean Shapley values, it does not extend naturally to higher-order moments. In particular, when applied to Gaussian processes, computing the variance or covariance structure of Shapley values is nontrivial, since orthogonality does not generally hold for product kernels. As a result, our method cannot be directly extended to explain GP models in a stochastic way.
By contrast, Mohammadi et al. [2025b] focus on FANOVA GPs, whose kernels admit a functional ANOVA decomposition. Compared to product kernels, this class of kernels is more restrictive and often harder to train in practice. Nonetheless, the additional structure provides a key advantage: it enables exact polynomial-time computation not only of the mean but also of the variance and covariance of stochastic Shapley values—something that product kernels cannot achieve. This distinction highlights the central difference between their method and ours.
Appendix J How does PKeX-Shapley differ from RKHS-SHAP [Chau et al., 2022]?
RKHS-SHAP is the first kernel method-specific SHAP-based algorithm. While the author still employs the conditional expectation value function ${\tilde{\nu}_{\boldsymbol{x}}}(S)=\mathbb{E}[f(X)\mid X_{S}=\boldsymbol{x}_{S}]$, they used the fact that for function $f$ in the RKHS $\mathcal{H}_{k}$, it is possible to estimate ${\tilde{\nu}_{\boldsymbol{x}}}(S)$ non-parametrically utilizing a tool known as conditional kernel mean embedding. Specifically, this leads to the following expression:
${\tilde{\nu}_{\boldsymbol{x}}}(\mathcal{S})=\boldsymbol{\alpha}^{\top}\mathbf{K}(\mathbf{K}_{\mathcal{S}}+n\lambda I)^{-1}k_{\mathcal{S}}(\mathbf{X}_{\mathcal{S}},\boldsymbol{x}_{\mathcal{S}})$
which of course, is different from our functional baseline value function
${\nu_{\boldsymbol{x}}}(\mathcal{S})=\boldsymbol{\alpha}^{\top}k_{\mathcal{S}}(\mathbf{X}_{\mathcal{S}},\boldsymbol{x}_{\mathcal{S}}).{}$
Although looking different in their expression, there is an interesting mathematical connection that is not apparent at first glance. First of all, we need to understand that conditional expectations can be interpreted as orthogonal projections to the space of nice-behaving functions defined on the conditioning variable. Specifically, let $\mathcal{L}^{2}(\mathcal{X})$ be the space of square-integrable functions on $X$, then for a random variable $Y$, the conditional expectation $\mathbb{E}[Y\mid X]$ is the unique element in $\mathcal{L}^{2}(\mathcal{X})$ that is closest to $Y$ in mean-square error:
$\mathbb{E}[Y\mid X]=\arg\min_{Z\in\mathcal{L}^{2}(\mathcal{X})}\mathbb{E}[(Y-Z)^{2}].$
As a result, this interpretation of conditional expectations, and thus of the conditional-expectation-based value function $\tilde{\nu}_{\boldsymbol{x}}$, allows us to directly connect it to our functional baseline value function. As established in Proposition 2 and Appendix B.3, the latter can also be viewed as an orthogonal projection, albeit onto a different function space.